# What Media Frames Reveal About Stance: A Dataset and Study about Memes in Climate Change Discourse

### v2

@@ 6 Meme Understanding through the Lens of Communication Science
To understand which aspects challenge models in meme understanding, we apply an interdisciplinary approach, integrating three concepts from communication science research: humor type, personalization, and responsibility. These concepts are critical in understanding the construction of meaning in affective climate communication on social media and may pose challenges for detection tasks. Humor is a key feature of memes as they are usually created with the intention to entertain people  Taecharungroj and Nueangjamnong (2015). Thus, we analyzed different humor types, such as puns, sarcasm, and surprise, which can have varying effects on readers. Personalization is a common communication strategy in political communication in general and that simplifies complex political issues by focusing on individual actors instead of groups.
We recruited two bachelor’s students in communication science to annotate humor type, person, and responsibility on 235 ClimateMemes test memes. Our guidelines are adapted from a comprehensive codebook on “Climate Change and Social Media” [^5] Anonymized version available at: https://osf.io/3hqdk?view_only=dd6035e7b03542e4a66c2fafa4bf0d7d allowing multiple labels per item (detailed definitions in Appendix G). Table 5 shows the most common labels, number of relevant memes, and LLaVA’s subsample performance.
In stance detection, LLaVA performs well on memes with humor types exaggeration and personification. Memes with irony and puns are difficult, receiving the lowest accuracy and F1. Memes with ordinary and celebrity under personalization categories show strong performance, but NGO stands out as challenging. For responsibility, memes concerning the macro-level are the easiest for the model, while meso-level memes are the hardest. Since micro/meso-level responsibilities address specific individuals (e.g., politicians) or organizations (e.g., companies), we hypothesize that macro-level responsibility (e.g., the society) leads to less variation and eases meme understanding.
@@
Concept Label #M Stance Frame
Acc. F1 Acc. F1
humor type irony 33 78.79 53.30 64.91 57.43
compare 25 88.00 69.90 65.75 54.67
surprise 21 85.71 50.42 64.58 56.82
personification 21 95.24 82.05 67.26 61.39
joke 19 84.21 52.53 58.14 50.86
exag 10 100.00 100.00 60.31 56.37
pun 5 80.00 33.33 59.69 42.86
Total 134 87.42 63.21 62.95 54.34
personalization ordinary 86 88.37 66.60 62.55 54.03
celebrity 25 88.00 53.23 61.75 52.27
politicial 14 85.71 78.79 54.80 47.93
NGO 14 50.00 37.18 66.74 55.43
Total 139 78.02 58.95 61.87 54.35
responsibility macro 50 88.00 51.06 64.88 58.39
meso 37 89.19 31.43 60.47 52.10
micro 37 83.78 46.88 60.26 52.56
Total 124 86.99 43.12 61.84 54.40
Table 5: Llava 4-shot meme+hum results on test subsamples with humor, person, and responsibility labels.
@@ 7 Conclusion
We introduce ClimateMemes, a new dataset of climate change memes annotated with stance and media frames. We demonstrate that media frame preferences are strong indicators of stance, with convinced and skeptical stances favoring distinct frames. We compare VLMs and LLMs and identify challenges in understanding multimodal information. Our interdisciplinary paper integrates concepts from communication science and reveals which meme aspects challenged the model.
@@
Our dataset is exclusively composed of memes collected from Reddit, which introduces a potential bias. By focusing solely on this platform, we limit the diversity of content that could be found on other platforms like Twitter, Facebook, Instagram, or 4chan. Each platform has its own user base, culture, and way of sharing and discussing memes, which could result in differences in the types of memes that are shared. This platform-specific limitation means that our findings might not be fully representative of meme trends across the internet as a whole. The memes on Reddit, for instance, might reflect particular political, social, or cultural preferences that do not necessarily align with those on other platforms, making the sample less generalizable.
Distribution of Meme Stances and Frames Might Be Different on Other Platforms
The stance and framing of memes on Reddit may not reflect the patterns found on other social media platforms. Reddit has a unique structure, where specific subreddits cater to distinct interests, communities, and ideologies, which could influence the stances and frames adopted in memes. For example, some subreddits may have a higher concentration of memes that are either supportive or skeptical of climate change, while other platforms might exhibit different trends. Memes on Twitter or Instagram could carry different connotations, tones, or styles that might not be as prevalent on Reddit. Thus, the distribution of meme stances and frames could vary significantly across platforms, and a more comprehensive understanding of meme discourse would require analyzing multiple platforms to account for these differences.
@@ Appendix C Frame Distribution in Climate Change Memes
Frame hoax cause impact allocation propriety adequacy prospect Total
Count 269 387 370 395 208 382 419 69 2,499
% 10.8% 15.5% 14.8% 15.8% 8.3% 15.3% 16.8% 2.8% 100%
@@ Appendix F Additional Experimental Results
Table 12: LLMs’ performance in terms of accuracy and Macro-F1 on stance detection. Hum = hum caption, syn = synthetic caption, #S = number of shots.
@@
(a) Stance detection
(b) Frame detection
@@ Appendix G Definition of CC-associated communication science concepts
For the humor types category, the content format used to create humor in memes is coded. Humor types are initially independent of the respective humor style. Following (Taecharungroj and Nueangjamnong, 2015), a distinction is made between seven humor types, several of which can in principle be used simultaneously in a meme.
@@
  • Comparisonsare combinations of two or more elements (e.g. before and after pictures) to construct a funny situa-tion.
  • Irony and sarcasm refers to the use of words to express the opposite of what one actually means.

### v3

@@ 6 Meme Understanding through the Lens of Communication Science
Concept Label #M Stance Frame
F1 Acc. F1 Acc.
humor
type
irony 33 53.30 78.79 57.43 64.91
compare 25 69.90 88.00 54.67 65.75
surprise 21 50.42 85.71 56.82 64.58
personif. 21 82.05 95.24 61.39 67.26
joke 19 52.53 84.21 50.86 58.14
exagger. 10 100.00 100.00 56.37 60.31
pun 5 33.33 80.00 42.86 59.69
[…]
personal
-ization
ordinary 86 66.60 88.37 54.03 62.55
celebrity 25 53.23 88.00 52.27 61.75
politicial 14 78.79 85.71 47.93 54.80
NGO 14 37.18 50.00 55.43 66.74
[…]
respons
-ibility
macro 50 51.06 88.00 58.39 64.88
meso 37 31.43 89.19 52.10 60.47
micro 37 46.88 83.78 52.56 60.26
[…]
Table 5: Llava 4-shot meme+hum results on test subsamples with humor, person, and responsibility labels. Best and worst scores per model are bolded and underlined.
To understand which aspects challenge models in meme understanding, we apply an interdisciplinary approach, integrating three concepts from communication science research: humor type, personalization, and responsibility. These concepts are critical in understanding the construction of meaning in affective climate communication on social media and may pose challenges for detection tasks. Humor is a key feature of memes as they are usually created with the intention to entertain people  Taecharungroj and Nueangjamnong (2015). Thus, we analyzed different humor types, such as puns, sarcasm, and surprise, which can have varying effects on readers. Personalization is a common communication strategy in political communication in general, and it simplifies complex political issues by focusing on individual actors instead of groups.
We recruited two bachelor’s students in communication science to annotate humor type, person, and responsibility on 235 ClimateMemes test memes. Our guidelines are adapted from a comprehensive codebook on “Climate Change and Social Media” [^5] https://osf.io/3hqdk?view_only=dd6035e7b03542e4a66c2fafa4bf0d7d provided by Luebke et al. (2025), allowing multiple labels per item (detailed definitions in Appendix G). Table 5 shows the most common labels, number of relevant memes, and LLaVA’s subsample performance.
In stance detection, LLaVA performs well on memes with humor types exaggeration and personification. Memes with pun and surprise are difficult, receiving the lowest F1 score. Memes with politicial and ordinary under personalization categories show strong performance, but NGO stands out as challenging. For responsibility, memes concerning the macro-level are the easiest for the model, while meso-level memes are the hardest. Since micro/meso-level responsibilities address specific individuals (e.g., politicians) or organizations (e.g., companies), we hypothesize that macro-level responsibility (e.g., the society) leads to less variation and eases meme understanding.
@@ 7 Conclusion
We introduce ClimateMemes, a new benchmark dataset of climate change memes annotated with stance and media frames. We demonstrate that media frame preferences are strong indicators of stance, with convinced and skeptical stances favoring distinct frames. We compare VLMs and LLMs and identify challenges in understanding multimodal information. Our paper also integrates concepts from communication science and reveals which meme aspects challenged the model.
@@
Our dataset is exclusively composed of memes collected from Reddit, which introduces a potential bias. By focusing solely on this platform, we limit the diversity of content that could be found on other platforms like Twitter, Facebook, Instagram, or 4chan. Each platform has its own user base, culture, and way of sharing and discussing memes, which could result in differences in the types of memes that are shared. This platform-specific limitation means that our findings might not be fully representative of meme trends across the internet as a whole. Therefore, the trends that we observed in our reddit meme data might not reflect trends beyond the period of investigation and the platform. Reddit has a unique structure, where specific subreddits cater to distinct interests, communities, and ideologies, which could influence the stances and frames adopted in memes. For example, some subreddits may have a higher concentration of memes that are either supportive or skeptical of climate change, while other platforms might exhibit different trends. Memes on Twitter or Instagram could carry different connotations, tones, or styles that might not be as prevalent on Reddit. Thus, the distribution of meme stances and frames could vary significantly across platforms, and a more comprehensive understanding of meme discourse would require analyzing multiple platforms to account for these differences.
Scope and Generalizability
This study focuses exclusively on climate change, a uniquely salient and persistently active topic with global relevance. Our analysis relies on a theoretical framework explicitly developed and validated for climate change discourse Jang and Hart (2015); Ross and Rivers (2019), examining context-specific media frames and stances, such as responsibility allocation (ALLOCATION) and the appropriateness of measures (PROPRIETY). As a result, our findings may not be directly generalizable to memes on other topics. Nevertheless, we believe that similar frames and patterns could apply to other topics, such as COVID-19 (Dynel, 2021), where the divide between convinced and skeptical stances exists alongside debates about responsibility and policy.
@@ Appendix C Frame Distribution in Climate Change Memes
Frame hoax cause impact
Count 269 387 370 395
% 10.8% 15.5% 14.8% 15.8%
Frame allocation propriety adequacy prospect
Count 208 382 419 69
% 8.3% 15.3% 16.8% 2.8%
@@ Appendix E Stance and Frame Distribution in Test Set & Majority Vote Baseline
Table 12: LLMs’ performance in terms of accuracy and Macro-F1 on stance detection. hum = human caption, syn = synthetic caption, #S = number of shots.
@@
Appendix F Additional Experimental Results
@@ Appendix G Definition of CC-associated communication science concepts
For the humor types category, the content format used to create humor in memes is coded. Humor types are initially independent of the respective humor style. Following Taecharungroj and Nueangjamnong (2015), a distinction is made between seven humor types, several of which can in principle be used simultaneously in a meme.
@@
  • Comparisons are combinations of two or more elements (e.g., before and after pictures) to construct a funny situation.
  • Irony and sarcasm refer to the use of words to express the opposite of what one actually means.